# Hyperparameters are all you need: Using five-step inference for an original diffusion model to generate images comparable to the latest distillation model

### v2

@@ 3.1 Algorithm Based on Truncation Error Analysis
$h_{t+1}=h_{t}+f_{\theta}(h_{t},t),$ (19)
$\frac{dh}{dt}=F_{\theta}(h_{t},t),$ (20)
where $h_{t}$ is the feature output in layer $t$.
The U-Net, as a black box ODE solver, can be regarded as another ODE. Thus, modifying the skip connection and scale backbone feature will further increase the norm of the output, increasing the moving distance of the diffusion reverse ODE. In the situation of the few-step inference, it will be a serious problems, since the weakness of the poor discretization method will be amplified by this behaviour. The related mathematic details will be mentioned below. We also visualize the decoded features along the trajectories of different samplers in fig.˜2.
Figure 2: The comparison between our custom DPM++1s sampler and the original DPM++1s sampler. The first line is the result trajectory of the original DPM++1s sampler with Karras’ schedule, the second one is Karras’ schedule + Free-U, and the last one is our sampler. The results show that if we apply Free-U in a few-step sampling without involving additional tricks, the generation result will be degraded.
To further show the influence of the Free-U on the black box ODE solver, we then visualize the decoded trajectory using different discrete methods by applying the same sampler DPM++ 1s sampler in fig.˜3. The standard time schedule using Free-U generates a trajectory similar to that of Karras et al. (2022) time schedule without Free-U, providing evidence that Free-U increases the moving distance at the beginning of the inference, which determines the image layout composition.
Figure 3: The comparison between different discrete methods when using the same DPM++ 1s Sampler in a few-step inference. The first line exhibits the trajectory of the normal Karras discrete method. The second one exhibits the trajectory of the Karras discrete method + Free-U. The third exhibits the normal discrete method with a normal U-Net, and the final one exhibits the normal discrete method with Free-U. And the final line is similar to the first line.
Another method is to parameterize the start step of using the U-Net decorator in an existing solver. However, experimental results show that this Vallina method doesn’t work either. We visualize the 6-step inference result combining the free-U and UniPC solver in fig.˜4.
@@
To address black-box ODE-solving problems, we should identify the characteristics of the diffusion ODE.
The basic logic of the ODE solver is to discretize a continuous curve trajectory into a series of lines. And to achieve the best performance, the ODE solver should determine the length of each discrete line for different ODEs, such as the Fehlberg method (see e.g. Atkinson et al. (2009)) that adjusts the discrete length based on the error per unit stepsize.
For the diffusion ODE, when the moving distance in step $t$ to $t+1$ is large during inference, the weakness of the poor discretion in the noisy stage will not be as fatal as the non-noisy steps. These insights have been proved by Karras et al. (2022) via an experiment that discretizes the inference process with a uniform gap between each timestep. And it obtains an array of steps $\{t_{i}\}$. Then, the researchers calculate the root mean square error (RMSE) between the one-step inference from $t_{i}$ to $t_{i+1}$ and the 200-step inference from $t_{i}$ to $t_{i+1}$. The experimental results show that the truncation error increases monotonically as the noise variance add in the input feature at the time step $t_{i}$ decreases.
[…]
Based on this observation, Karras et al. (2022) proposes a following discrete method:
eq.˜21 controls the variant $\sigma_{i}$ via $p$, and if the $p$ increases, the $\sigma_{i}-\sigma_{i-1}$ will increase in the noisier part and decrease in the less noisy part. Moreover, according to the observation from the Li & Zhang (2025), if we choose $p=7$ and $N=20$, the fourth-to-last step corresponds to the first $\frac{6}{1000}$ perturbation steps, and it will cost $\frac{1}{5}$ inference budget. Considering the U-Net has 860 million parameters, and the variance of the $p(x_{t}|x_{0})$ is approximately $\frac{6}{1000}$, performing four-step ODE inference at this point is illogical. Moreover, the $\beta$-VAE used for decoding is also trained with a small KL-term constraint, which enables it to handle the noise in the latent variable.
Based on that observation, Li & Zhang (2025) proposes another discrete method:
When choosing $p=1.2$, the $h_{i}$ in both the noisy part and the less noisy part are larger than those in Karras’s original method. In such a situation, $t_{i}-t_{i-1}$ will increase in the noisier part. The eq.˜4 shows that $t_{i}$ is proportional to the variance of noise added in $x_{t}$, which can be deemed as the speed of movement at timestep $t_{i}$. The increase of $t_{i}-t_{i-1}$ in the noisy part aligns with the rise in speed, while the rise of $\sigma_{i}-\sigma_{i-1}$ aligns with the increment of the moving distance, hence it can cope with the problem of the original Karras schedule mentioned above.
However, both of those discrete methods have been commonly used in all kinds of images. Our research shows that, even with the same time schedule and UniPC solver, perceptual quality in the few-step inference remains substantially different across image resolution and image type. Hence, the design space of hyperparameters for different ODE solvers has not yet been totally exploited. We find a discrete method, which changes the eq.˜23, a formula with a special design reason mentioned later, by adding one more step to get clean denoised images. And use it to generate images at different resolutions and of different types in a few-step inference. We find that the result in fig.˜5 has a noticeable gap in perceptual quality.
Figure 5: The comparison between the different resolution images and different type images. The first line contains 512 x 512 images generated by Stable Diffusion 1.5, the second line contains 1024 x 1024 images generated by Dreamshaper XL. The last lines contain high-resolution ACGN images generated by Illustrij Evo.
Based on the experimental results, designing a customized training-free algorithm for different image types requires exploring the design space of the ODE solver for the latent diffusion model.
@@
Firstly, the decoder of the latent diffusion model, a $\beta$-VAE trained with a small KL-term constraint, can handle a small amount of noise; thereby, if we choose a time schedule that does not fully denoise the image, we can save further inference budget. Although the latent diffusion model theoretically does not require using the KL term to train an autoencoder, this optimization trick is very prevalent in the training of the diffusion model. The latest representation autoencoders (RAE) also add noise during the training to augment the decoder Zheng et al. (2025). The experimental results show that the decoder’s capacity to handle noisy features varies with image resolution. In the fig.˜6, the second line uses the same time schedule as the first line, which doesn’t denoise totally, but those images in the second line have some blurred parts that harm the details of the images, while the first line doesn’t confront this problem. And the third line output feature with less noise, and has a better results.
@@
where $\sigma_{stop}$ is defined via the following formula:
@@
which could be nonzero to save the one-step inference budget. By applying this special time schedule, the problems of using Free-U while preventing an increase in moving distance could be addressed, since we could only apply Free-U at timesteps that move a short distance. By leveraging the robustness of the $\beta$-VAE, the total moving distance has also been constrained.
The visual difference between our discrete method and the method proposed by the Karras’ paper is in fig.˜7
@@
Meanwhile, since the total moving distance should be as limited as possible, we also apply the analytical first step (AFS) proposed by Dockhorn et al. (2022) during inference.
Our algorithm can be described by the algorithm˜1
@@
1 $\sigma_{stop}=(\sigma_{max}^{\frac{1}{p_{2}}}+\frac{t_{stop}}{N+t_{stop}+1}(\sigma_{min}^{\frac{1}{p_{2}}}-\sigma_{max}^{\frac{1}{p_{2}}}))^{p_{2}}$;
2
[…]
4if afsafs then
@@
14
15 end while
Algorithm 1 custom Karras scheduler in the common framework of Diffusion ODE solving
@@ 3.2 Information Theory Analysis of the Different Algorithms
The experimental results in the section˜4 show that the diversity of the generative images sampled from a distillation model is the main bottleneck of its FID performance. Given that all images are generated from the same prompts, and the distillation algorithm changes the score function from pointing to the centroid of all possible images to a special image, we assume the distillation model limits the diversity of images’ low-frequency components, which is caused by modification of the score function at the start stage of the inference. Explicitly, some images with totally different semantics could have a similar, fuzzy structure with the same color block in a similar position. And the limitation on the possible fuzzy structure of images may also cause diversity problems. And we use information theory to explain why applying Free-U only in the final stage of the inference helps address this problem.
To introduce the relationship between information theory and machine learning, we should first review the historical confusion about machine learning’s generalization capability.
A common point of confusion in deep learning at an early stage is how neural networks with so many parameters can still generalize (Breiman (2018)). The standard statistical model uses a limited number of parameters to concisely describe the phenomenon and capture the essence of the observed data, whereas deep learning is fundamentally different. For this problem, the study theory (Harman & Kulkarni (2012)) claims that limiting the number of functions that a model can express helps ensure that the model approaches the best functions it can have, and that it also bounds the distance between the final function a neural network study about and the average of functions.
Generally, a possible function expressed by a model is a hypothesis $\mathcal{H}$, and if the number of input data points is $N$, the number of possible hypotheses is less than $2^{N}$. According to information theory, for a dataset $\mathcal{S}$ with data $D$, the typical set that a model can access is $2^{H(D)}$ (Yeung (2012)). For example, in a high-dimensional standard Gaussian distribution, a pure black image, although with the highest possible density, is not in the typical set. Considering the amount of data in the typical set, and the mutual information between the output and the middle feature of the neural network $I(y,z)$, we can further bound the models’ capacity for potential hypotheses.
Explicitly, we can consider the probability of middle feature $z$ given output $y$, and the $p(z|y)$ provides a match between $y$ and $z$. The information bottleneck theory proposed by Tishby et al. (2000) shows that $2^{H(Y|Z)}$ denotes the volume of $Y$ projected to the $Z$. Hence, $2^{I(Y,Z)}=\frac{2^{H(Y)}}{2^{H(Y|Z)}}$ further constraints the possible hypotheses by limiting the possible middle feature (Tishby et al. (2000)). As the mutual information bounds the distance between the study hypothesis and the average hypothesis, generative adversarial networks (GANs) (Goodfellow et al. (2020)), which face mode collapse (Creswell et al. (2018)) that only generate a subset of the dataset, can benefit from it by adding a mutual information constraint Jeon et al. (2021). One study (Li & Zhang (2025)) shows that this phenomenon also appears in the diffusion distillation, which consistently changes the score function and triggers an increase in mutual information. The basic idea of the proof in that study is related to $\beta$-variation auto-encoder ($\beta$-VAE). When $\beta$-VAE decreases $\beta$ to let VAE generate high-quality images, the upper bound of the mutual information constraint will be abated. Meanwhile, the peak signal-to-noise ratio (PSNR) of the encoding feature $Z$ will increase. From this perspective, the diffusion model can be deemed as a hierarchy VAE (Luo (2022)), while progressive distillation proposed by Salimans & Ho (2022) increases the mutual information by letting the model directly use $X_{\theta}({t+2})$ to replace the original output $X_{\theta}(t+1)$. Where $X_{\theta}$ is a noisy feature output by solving ODE. Hence, the PSNR of the middle output feature also increases.
@@
However, it is intuitive that changing the direction of the score function to allow it to point to the final result, rather than the centroid of the possible image, will also affect the mutual information, even if the output does not have an increased PSNR. We provide further analysis regarding this.
The diffusion ODE can be regarded as a Markov process $Z_{\theta}(T)-Z_{\theta}(T-1)-...-Z_{\theta}(t)-...Z_{\theta}(0)$, in which $Z_{\theta}(t)$ is the noisy input of the diffusion model, and we have the following eq.˜25:
$I(Z_{\theta}(0),Z_{\theta}(t))=H(Z_{\theta}(0))-H(Z_{\theta}(0)|Z_{\theta}(t))$
$\leq H(X(0))-E_{Z_{\theta}(t)}[p(Z_{\theta}(0)|Z_{\theta}(t))$
$*\log p_{\theta}(Z_{\theta}(0)|Z_{\theta}(t)))],$ (25)
in which $H(X(0))$ is the entropy of the real $X(0)$ sample from the dataset.
In $\beta$-VAE, when given the random variable $Z$, the output $Y$ is assumed to be sampled from a Gaussian distribution using $Y$ as its mean, and abating the $\beta$ makes the middle feature $Z$ more similar to $Y$, enhancing the log likelihood of the $Y$. And in the distillation algorithm, eq.˜25 shows that if we modify the $Z_{\theta}(t)$ to make it similar to the final output to enhance the log-likelihood by rectifying the trajectory, then the upper bound of the mutual information will also increase. Most distillation methods that do not increase the PSNR of the output will still face this problem. We evaluate the Precision and Recall of Distribution (PRD) (Sajjadi et al. (2018)) of flash diffusion and our sampler to further validate this claim.
@@ 4.1 Ablation Study
We conduct an ablation Study on a customized UniPC solver that aims to generate a 1024 x 1024 image in a few-step inference guided by 5000 prompts sampling from the COCO 2017 dataset. Explicitly, for the 8-step inference, we employ $p_{1}=12$, $p_{2}=1.2$, $stop=10$, $M=12$, and $N=8$ as the discrete setting. And for the 6-step inference, the discrete setting is $p_{1}=12$, $p_{2}=1.2$, $stop=10$, $M=12$, and $N=6$. Then, we try the following setting: 1. apply Free-U on every step of inference, 2. apply Free-U only on the last few steps of inference, 3. Don’t apply Free-U. The experimental result is in the table˜1. The $w$ in the table˜1 is the guidance scale of inference, and the number in the parentheses is the total number of function evaluations (NFEs) of the model. Meanwhile, for the 8-step UniPC solver, we use a 2nd-order multistep algorithm, whereas the 6-step solver is 1st-order solver.
Solver FID CLIP-Score
custom UniPC(6) apply Free-U on the every steps $w=5.5$ 28.88 30.96
custom UniPC(8) apply Free-U on the every steps $w=5.5$ 33.70 30.50
custom UniPC(6) don’t apply Free-U $w=5.5$ 27.13 31.49
custom UniPC(8) don’t apply Free-U $w=5.5$ 25.47 31.39
custom UniPC(6) start apply Free-U at 4th step $w=5.5$ 25.20 31.42
custom UniPC(8) start apply Free-U at 6th step $w=5.5$ 24.38 31.38
custom UniPC(6) apply Free-U on the every steps $w=7.5$ 39.62 30.33
custom UniPC(8) apply Free-U on the every steps $w=7.5$ 54.26 29.63
custom UniPC(6) don’t apply Free-U $w=7.5$ 25.34 31.53
custom UniPC(8) don’t apply Free-U $w=7.5$ 24.61 31.22
custom UniPC(6) start apply Free-U at 4th step $w=7.5$ 25.29 31.44
custom UniPC(8) start apply Free-U at 6th step $w=7.5$ 24.58 31.21

### v3

@@ 3.1 The effect of the skip-connection modification
$l_{t+1}=l_{t}+f_{\theta}(l_{t},t),$ (8)
$\frac{dl}{dt}=F_{\theta}(l_{t},t),$ (9)
where $l_{t}$ is the feature output in layer $t$.
The U-Net, as a black box ODE solver using skip connections, can be regarded as another ODE. And modifying the skip connection and scale backbone feature in the neural ODE will further increase the $h_{i}$ in Equation 4.
To further show the influence of the Free-U on the black box ODE solver, we then visualize the decoded trajectory using different discrete methods by utilizing the same sampler DPM++ 1s sampler in Figure 3. The uniform time schedule using Free-U generates a feature in the first two steps that is similar to the Karras et al. (2022) time schedule without Free-U, while totally different to the normal uniform sampler without using Free-U, providing evidence that Free-U increases the moving distance at the beginning of the inference. And if we use karras or a similar time schedule, it will have an unwilling effect, since it’s large enough at the beginning step.
Figure 3: The comparison between different discrete methods when using the same DPM++ 1s Sampler in a few-step inference. The first line exhibits the trajectory of the normal Karras discrete method. The second one exhibits the trajectory of the Karras discrete method + Free-U, which has an unwilling result. The third exhibits the normal discrete method with a normal U-Net, and the final one exhibits the normal discrete method with Free-U. And the final line is similar to the first line.
To prevent this problem, another method is to parameterize the start step of using the U-Net decorator in an existing solver. However, experimental results show that this Vallina method doesn’t work either. We visualize the 6-step inference result combining the free-U and UniPC solver in Figure 4.
@@
To address black-box ODE-solving problems, we should provide a new time schedule.
3.2 The time schedule leveraging $\beta$-VAE
@@
The decoder of the latent diffusion model, a $\beta$-VAE trained with a small KL-term constraint, can handle a small amount of noise; thereby, if we choose a time schedule that does not fully denoise the image, we can save further inference budget. Although the latent diffusion model theoretically does not require using the KL term to train an autoencoder, this optimization trick is very prevalent in the training of the diffusion model. The latest representation autoencoders (RAE) also add noise during the training to augment the decoder (Zheng et al., 2025). The experimental results show that the decoder’s capacity to handle noisy features varies with image resolution. In the Figure 5, the second line uses the same time schedule as the first line, which doesn’t denoise totally, but those images in the second line have some blurred parts that harm the details of the images, while the first line doesn’t confront this problem. And the third line output feature has less noise and produces better results.
@@
where $\sigma_{stop}$ is defined via Equation 11, and $t(\sigma)$ is a function projecting variant $\sigma$ to timestep $t$, which use the discrete version describe in Equation 1 whose range is between $[1,1000]$:
@@
and $\sigma$ could be nonzero to save the one-step inference budget by utilizing $\beta$-VAE. Meanwhile, the problems of using Free-U while preventing an increase in moving distance could be addressed by applying this special time schedule, since we could only apply Free-U at timesteps that move a short distance. By leveraging the robustness of the $\beta$-VAE, the total moving distance has also been constrained. The visual difference between our discrete method applying on DPM++ 2m, and the method proposed by the Karras’ paper is in Figure 6
@@
Meanwhile, since the total moving distance should be as limited as possible, we also apply the analytical first step (AFS) proposed by Dockhorn et al. (2022) when generating high-resolution images.
Our algorithm can be described by the Algorithm 1
@@
$\sigma_{stop}=(\sigma_{max}^{\frac{1}{p_{2}}}+\frac{t_{stop}}{N+t_{stop}+1}(\sigma_{min}^{\frac{1}{p_{2}}}-\sigma_{max}^{\frac{1}{p_{2}}}))^{p_{2}}$
$ts=\{t_{i}=(t_{max}^{\frac{1}{p_{1}}}+\frac{i}{N}(t(\sigma_{stop})^{\frac{1}{p_{1}}}-t_{max}^{\frac{1}{p_{1}}}))^{p_{1}};0\leq i<N\}$
if afsafs then
@@
8
9end while
Algorithm 1 F-scheduler in the common framework of Diffusion ODE solving
@@ 3.3 Information Theory Analysis of the Different Algorithms
The experimental results in the Section 4 show that the diversity of the generative images sampled from a distillation model is the main bottleneck of its FID performance. To explain why our algorithm can outperform the distillation model, we use information theory to provide a tentative analysis of the diffusion model.
Bertrand et al. (2025) shows that using the ground truth score function in the place having the highest loss, explicitly, 20% of inference, to replace the studied score function will harm the diversity of the diffusion model. They don’t provide a theoretical analysis of this phenomenon, nor do they provide a further estimate of the distillation model that modifies the score function. In this section, the information bottleneck theory Theorem 3.1 reveals that mutual information affects the model’s generalization. We find out when the score function been changed from the centroid of all possible images that can generate the current noisy image to a special image, the mutual information between the middle feature and the output increases, as described in Proposition 3.2. By preventing the mutual information from affecting, Theorem 3.1 and Proposition 3.3 show that the current average output of the model would not shift and deviate from the ground truth centroid, while the mode collapse in GAN, shown in Figure 7, is different from that. Jeon et al. (2021) also shows that constraining the information bottleneck in GAN can enhance the FID.
Eplicilty, for an arbitrary image generative model, we assume that there is a ground truth noise $x$ and image $y$ pair that is possible to be implicitly studied by the neural network, and $y$ sample from $\mathcal{Y}$ contains images with enough diversity.
Then, according to learning theory, if we want to prevent a model from rote, we should constrain the number of functions this model can express, to ensure it can get access to the best hypothesis in training.
Suppose we have a training data $s_{n}=\{y_{i}\}_{i=1}^{n}$ that has at least $1-\sigma$ probability drawn from the same distribution of $\mathcal{Y}$, a neural network $f^{s}$ trained in $s$ via SGD in $i$ iteration, a batch size $m$ and a batch validating data $s_{m}=\{\hat{y}_{i}\}_{i=1}^{m}$, a middle feature $Z_{l}^{s}$ output from the layer $l$ of the neural network $f^{s}$. Then, for the output of the $f^{s}$, we consider a output $f^{s}(\hat{x}_{1}),f^{s}(\hat{x}_{2}),...,f^{s}(\hat{x}_{m})]$ from a network on $i$-th iteration traning which the $m$ noise-image pair $\hat{x},\hat{y}$ is denote by the ground truth matching. We consider the finite hypothesis set $\mathcal{H}$ that consists of all possible outputs in the $i$-th iteration of SGD training, without considering the order of outputs within the same batch. Meanwhile, in SGD, the training process is random, and with a given noise-feature pair $x,z_{l}^{s}$ in training, we assume $p(Z_{l}^{s}|Y)$, who denote the probability density of $Z_{l}^{s}$ in the $i$ iteration of SGD training, will use $z_{l}^{s}$ as the means. And the mutual information $I(Y;Z_{l}^{s})$ is given by $p(Z_{l}^{s},Y)$. Finally, seperate the neural network into two parts: decoder $g^{s}$ and encoder $\phi^{s}$, and $f^{s}=g^{s}\circ\phi^{s}$, where we use $\circ$ to denote the composition of functions.
First, the following theorem investigates the information bottleneck (IB) of the latent diffusion model, whose proof is listed in the Appendix A.
Theorem 3.1.
For the L2 loss function $\mathcal{L}$, and $\Delta(s)=E_{X,Y}[\mathcal{L}(f^{s}(X),Y)]-\frac{1}{m}\Sigma_{i=1}^{m}\mathcal{L}(f^{s}(x_{i}),y_{i})$, we have:
$\Delta(s)\leq\underset{l}{\arg\min}\sqrt{\frac{2^{mI(Y;Z_{l}^{s})}+{I(\phi_{l}^{s};s_{n})}+log\frac{2}{\sigma}}{2o*A_{m}^{m}}},$ (12)
where $o$ is a fixed constant for every hypothesis space, and $A_{m}^{m}$ denotes the number of arrangements of $m$ distinct objects.
Unlike (Kawaguchi et al., 2023) that applies IB to a stochastic neural network, we also show that, even without the stochastic part in the neural network, the SGD itself will give the mutual information between the middle feature and the input, and it still constrains the possible hypotheses. Add a stochastic process to the neural network, or use the loss with high variance, just a special case of our analysis.
@@
Then, it is intuitive that changing the direction of the score function to allow it to point to the final result, rather than the centroid of the possible image, will affect the mutual information, then affect the Equation 19, and cause the diversity problem. We provide further analysis regarding this. Explicitly, the diffusion ODE can be regarded as a Markov process $Z_{\theta}(T)-Z_{\theta}(T-1)-...-Z_{\theta}(t)-...Z_{\theta}(0)$, in which $Z_{\theta}(t)$ is the noisy input of the diffusion model, and we have the following Proposition 3.2:
Proposition 3.2.
$I(Z_{\theta}(0),Z_{\theta}(t))=H(Z_{\theta}(0))-H(Z_{\theta}(0)|Z_{\theta}(t))\leq H(Z(0))-E_{Z_{\theta}(t)}[-p(Z_{\theta}(0)|Z_{\theta}(t))*\log p_{\theta}(Z_{\theta}(0)|Z_{\theta}(t)))]$
in which $H(Z(0))$ is the entropy of the real $Z(0)$ sample from the dataset. If we modify the $Z_{\theta}(t)$ to make it similar to the final output to enhance the log-likelihood by rectifying the trajectory, then the upper bound of the mutual information will also increase.
In the following proposition, we decompose the $E_{X,Y}[\mathcal{L}]=E_{X,Y}[\mathcal{L}(f^{s}(X),Y)]$ into two parts to explain mode collapse. And $\Delta(s)$ been constrain by IB above is the uppper bound of the $E_{X,Y}[\mathcal{L}]$
Proposition 3.3.
$E_{X}[(f^{s}(X)-\hat{f}(X))^{2}]=E_{X,Y}[\mathcal{L}]-E_{X,Y}[(\hat{f}(X)-y)^{2}]$, where $\hat{f}(x)=E_{s_{n}}[f^{s}(x)]$
Meanwhile, as the Proposition 3.2 explain, the constrain of mutual information also reduce the log likelihood related to the second term. And, in GAN, the mode collapse will enlarge the $E_{X}[(f^{s}(X)-\hat{f}(X))^{2}]$, like the Figure 7. Hence Proposition 3.3 can prevent the diversity problem. The proof of Proposition 3.3 is postponed to Appendix B.
Figure 7: Illustration of Mode Collapse. (Left) The target distribution consists of a mixture of 8 Gaussians arranged in a ring. (Right) A simulated example of mode collapse that contains only a subset of the available modes.
@@ 4.1 Ablation Study
To demonstrate the necessity of applying Free-U at different steps, we conduct an ablation Study on a customized UniPC solver that aims to generate a 1024 x 1024 image in a few-step inference guided by 5000 prompts sampled from the COCO 2017 dataset. Explicitly, for the 8-step inference, we employ $p_{1}=12$, $p_{2}=1.2$, $stop=10$, $M=12$, and $N=8$ as the discrete setting. And for the 6-step inference, the discrete setting is $p_{1}=12$, $p_{2}=1.2$, $stop=10$, $M=12$, and $N=6$. Then, we try the following setting: 1. apply Free-U on every step of inference, 2. apply Free-U only on the last few steps of inference, 3. Don’t apply Free-U. The experimental result is in the Table 1. The $w$ in the Table 1 is the guidance scale of inference, and the number in the parentheses is the total number of function evaluations (NFEs) of the model. Meanwhile, for the 8-step UniPC solver, we use a 2nd-order multistep algorithm, whereas the 6-step solver is 1st-order solver.
@@
Solver steps use Free-U apply stage FID CLIP-Score
UniPC 6 Yes Every Steps 28.88 30.96
UniPC 8 Yes Every Steps 33.70 30.5
UniPC 6 No Never 27.13 31.49
UniPC 8 No Never 25.47 31.39
UniPC 6 Yes 4th step 25.20 31.42
UniPC 8 Yes 6th step 24.38 31.38